OpenStudy (samsan9):

what is the vertex of the graph of this function y=-(x+2)(x+4)? A.) (3,-35) B.) (-3,1) C.) (-2,-4) D.) (-3,-1)

OpenStudy (imstuck):

The x intercepts are given to you within your sets of parenthesis. x + 2 = 0...x = -2. x + 4 = 0...x = -4. So the points where the graph crosses the x axis are at -2 and -4. The x coordinate of the vertex is at x = -3, because the vertex is centered between the two x intercepts. Let's find y now...give me a sec to dwell on that one...

OpenStudy (imstuck):

You have to complete the square on this to find the y coordinate. That works out like this:

OpenStudy (samsan9):

do you mean (x+2)(x+4) because i got -x^2-6x-8

OpenStudy (imstuck):

Hold on, let me finish something first ok? Don't multiply the negative into the polynomial; you just have to take it out again. so leave it out in front. It makes things easier.

OpenStudy (tkhunny):

x = -3 y=-((-3)+2)((-3)+4) = -(-1)(1) = 1 (-3,1)

OpenStudy (imstuck):